OpenStudy (anonymous):

Find the slope of the line containing the following two points: (3/10,-3) and (-1/5,-1/5) (BEST ANSWER GETS MEDAL!)

OpenStudy (anonymous):

Please please help! :)

OpenStudy (anonymous):

Slope would be 3.2/1/2 or 3 1/5 / 1/2

OpenStudy (anonymous):

so the answer is 3.2/0.5?

OpenStudy (anonymous):

yes mam.

OpenStudy (anonymous):

the system marked it wrong... weird

OpenStudy (anonymous):

hmm did it give an option in fraction form?

OpenStudy (anonymous):

Slope should be - 28 / 5

OpenStudy (anonymous):

Yay! That's right :D

OpenStudy (anonymous):

hmm show how u got that.. I think I need a refresher.

OpenStudy (anonymous):

- 1/5 + 3 14/5 ----------- = --------- = -140 / 25 = -28/5 -1/5 - 3/ 10 -5/10

OpenStudy (anonymous):

oh thanks skipped the second step and threw my math off. Appreciate it.

OpenStudy (anonymous):

Lol